OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points 7,5 and -4, 1.

OpenStudy (anonymous):

1. Which is an equation in point-slope form of the line that passes through the points 7,1 and -4,-1? A. y +1 = 6/7(x+4) B. y+ 1= 6/11(x+4) C. y+4= 11/6(x+1 D. y-1= 6/11(x-4)

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

@miah872

OpenStudy (anonymous):

@nbandy0128

OpenStudy (jordanloveangel):

ayeeee 1 sec

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

I will give medals to whoever gets the answer right. But first I have more questions after this one. When I submit my assessment and the answers are right, then ___________ will get a medal.

OpenStudy (anonymous):

And a fan.

OpenStudy (jordanloveangel):

1.c?

OpenStudy (anonymous):

You think so?

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

yes im not sure

OpenStudy (anonymous):

Ok, I'll put it.

OpenStudy (anonymous):

Which is an equation of the given line in standard form? A. –8x + 9y = 23 B. –8x + 9y = –23 C. –8x + 7y = 25 D. –9x + 8y = –23 http://static.k12.com/calms_media/media/1331000_1331500/1331297/1/c5c8e45fd70b35572ab9141472538d3ecb31705d/HS_AL1G2_S1_07_10_ART2.jpg

OpenStudy (anonymous):

Sorry with all the question mark my computer is messed up.

OpenStudy (anonymous):

The link to the diagram is in the question.

OpenStudy (jordanloveangel):

–8x + 9y = 23

OpenStudy (anonymous):

So it's A?

OpenStudy (jordanloveangel):

yeah

OpenStudy (anonymous):

Ok thanks.

OpenStudy (anonymous):

Which is the equation of the given line? A. y = –2x B. x = 2 C. x = –2 D. y = –2 http://static.k12.com/calms_media/media/308500_309000/308807/1/766e8c16127d3afa586e7a118386416b180b22d2/MS_ALG_S1_07_10_quiz_Q3_question.gif

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

@love10129151

OpenStudy (anonymous):

PLEASE HELP MY THIRD QUESTION

OpenStudy (anonymous):

@adamanderson

OpenStudy (jordanloveangel):

http://openstudy.com/study#/updates/50a1447be4b0e22d17eef601 @maddiemsp